# Remote Cooling of Spin-ensembles Coupled to a Spin-mechanical Hybrid Interface

### v1

@@ IV Probe spin control
[…]
Figure 2: Quantum circuit illustrating the implementation of a projector on the oscillator, influenced by the spin ensemble states and the measurement result of the probe spin. The process starts with the ancilla spin in the $\ket{+}$ state, followed by the application of repeated DD pulses. This leads to a $\mathcal{D}^{f}_{\scriptscriptstyle k,\pm}(t)$ controlled evolution guided by the spin ensemble $P_{k}$. The procedure ends with a $\pi/2$ pulse and measurement of the ancilla spin in the Pauli Z basis. The specific projector realized is expressed mathematically in Eq (15).
In our investigation, we focus specifically on the weak dispersive coupling between each spin and the oscillator, deliberately ignoring any direct interactions between the spins. Nevertheless, by carefully choosing the detuning $\epsilon$ and the number of pulses $n$, substantial changes in the dynamics of the spin ensemble and the oscillator can be induced. The modification of the dynamics is facilitated by a cyclic process involving the projective measurement and reinitialization of the ancilla spin, as schematically depicted in Fig. 1. Initially, the ancilla spin is prepared in the state $\ket{+}$. Then the system undergoes a DD-modulated evolution for a time $t=n\tau$, as described in Eq (3). Finally, the probe spin is measured in the Pauli-Z basis, following the application of an additional $\pi/2$ pulse to it.
The process described above is represented by the circuit in Fig. 2. It leads to the implementation of a projector throughout the system depending on the probe measurement outcome, as mathematically formulated:
$\mathcal{P}_{\scriptscriptstyle N,\pm}=\sum_{\scriptscriptstyle k=0}^{\scriptscriptstyle N}V_{k,\pm}\otimes\mathrm{I}_{k},\quad V_{k,\pm}=\frac{1}{2}\left(\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,+}(t)\pm\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,-}(t)\right),$ (15)
where $V_{k,\pm}$ is the projector on the oscillator, $k$ is the number of ensemble spins pointing up, and $\pm$ is the result of the ancilla measurement. The repeated application of such projectors is central to the cooling of the oscillator in our previous work [58], which is now extended to cooling/polarizing an entire spin ensemble in this work.
Figure 3: We plot the probability of projecting a spin ensemble into its possible total magnetization states (from $-N/2$ to $N/2$) after repeated conditional measurements of the probe spin in the state |$\ket{0}$, as described in Eq. (18). Here we consider an ensemble with $N=14$ spins, an oscillator with frequency $\omega=1.2$MHz that is weakly coupled to the probe and spin ensembles with strengths of $g_{0}=g=10$kHz. The probe is modulated periodically with $n=128$ $\pi$-pulses with an interval $\tau=\pi/{\omega-\epsilon}$. Here the detuning is set as $\epsilon=1.1$kHz. The probability distribution reaches saturation after $M=100$ measurements.
V System cooling
In this section, we describe the protocol that imposes a thermal filter on the hybrid system, allowing for controlled cooling of either the oscillator, the spin ensemble, or both. We start with $N$ spins in the ensemble, which are initially assumed to be in a fully mixed state, and the oscillator is assumed to be in its thermal state. These initial states are written as
[…]
Following the circuit depicted in Fig. 2, we measure the probe spin, selecting only outcomes of 0. This process effectively implements the projector $\mathcal{P}_{\scriptscriptstyle N,+}$ as defined in Eq. (15). Upon performing $M$ such measurements, the probability that the ensemble is in the $\mathrm{I}_{k}$ sector is calculated as
$\mathrm{P}_{m}=\frac{1}{\sum_{m}\mathrm{P}_{m}}\mathrm{Tr}[\mathcal{P}_{\scriptscriptstyle N,+}^{M}\,\rho_{B}(0)\,\mathcal{P}_{\scriptscriptstyle N,+}^{M}],$ (18)
where $m=(2k-N)/2$ is the total magnetization of the ensemble. This dependence on $m$ shows that, with appropriate parameter choices, it is possible to achieve $\mathrm{P}{\scriptscriptstyle m=N/2}\approx 1$, thereby polarizing the ensemble. This ensemble polarization depends non-trivially on the filter function imposed by the pulses on the probe spin, as can be seen from Eq. (14). It gives us a new control parameter $\epsilon$ that would affect the projection of the spin ensemble onto different $m$-states.
In this work, we will focus on the weak coupling regime, i.e., $g\ll\omega$, and we choose the detuning $\epsilon\sim\mathcal{O}(g)$ [31] so that $\epsilon\ll\omega$. Specifically, we consider the case where the spin couplings $g,~{}g_{0}$ are $10$ kHz and the mechanical oscillation frequency is chosen to be $1.2$ MHz. Consequently, for a spin ensemble of size $N=14$, by fixing the detuning to be $\epsilon=1.16$ kHz, one can see that the sector with all spins up ($m=7$) is predominantly left, effectively polarizing the ensemble. To further illustrate the non-trivial dependence of the maximum ensemble polarization on the detuning $\epsilon$, we vary its value from 0 to 2 kHz and also vary the number of ensemble spins. Note that we always assume $M=100$ projections of the probe spin.
Figure 4: Non-trivial dependence of the final maximum polarization on the detuning $\epsilon$ for varying spin-ensemble sizes ($N$). The y-axis represents the probability of finding all ensemble spins pointing up after $M=100$ probe projection measurements, denoted as $\mathrm{P}_{\scriptscriptstyle m=N/2}(M)$. Note that the oscillator frequency is set as $\omega=1.2$MHz and the spin-oscillator coupling strengths are set as $g_{0}=g=10$kHz.
A critical factor in achieving ensemble polarization is to ensure that the interval between successive pulses on the probe is close to resonance, i.e., a negligible detuning, $\epsilon\approx 0$, as indicated in Eq. (13). The reason for this is to prevent premature cooling of the oscillator, since its thermal occupancy significantly affects the interaction between the probe and the ensemble, as detailed in Eq. (1). To allow the oscillator to cool, the detuning must be set to a higher value to facilitate its rapid cooling [58].
In Fig. 5 we show a pulse sequence that polarizes an ensemble of $N=4$ spins and cools an oscillator starting from an initial thermal occupancy of $n_{0}=45$. We first optimize the cooling of the spin-ensemble by varying the detuning, to find an optimal point $\epsilon_{1}=1.1$kHz where the ensemble polarizes perfectly, while the oscillator occupancy remains unchanged. This behavior is similar to what one would expect for resonant case ($\epsilon=0$). Now for another $50$ probe measurements and shifting the detuning to a higher value of $\epsilon_{2}=5$kHz, would rapidly cool the oscillator as shown in the right part of Fig. 5. Shifting $\epsilon_{2}$ to even higher values would require more inversion pulses on the probe to achieve a similar effect, as can be seen from Eq. (12). We would like to note that there would be no effect of the dephasing caused by the surrounding spin-bath on the cooling dynamics, as each measurement also partially collapses the spin-bath towards steady states that eventually stabilize the cooling mechanism [60]. Other noise sources only affect the cooling rates, but not the final steady states of the ensemble and the oscillator.
@@ V System cooling
Figure 5: Cooling the oscillator and spin ensemble. This figure shows the polarization $\mathrm{P}_{z}$ and the thermal occupancy $n(\omega)$ for both the spin ensemble and the oscillator, illustrating how the cooling/polarization progresses with the number of measurements $M$. The effective cooling of these quantum systems occurs in different parameter regimes. Initially, the spin ensemble is cooled by adjusting the $\pi$ pulse intervals to near-resonance conditions, specifically $\tau_{1}$ such that $\epsilon_{1}=0.11$kHz, where the thermal occupancy of the oscillator remains largely unaffected. After the spin ensemble has cooled, changing the pulse interval to $\tau_{2}$, which produces a more significant off-resonance with $\epsilon_{2}=1.5$kHz, leads to a rapid cooling of the oscillator to its ground state. The coupling parameters are also $g_{0}=g=10$kHz, with an oscillator frequency $\omega=1.2$MHz. Here, we have considered an initial oscillator thermal occupation of $n_{0}\approx 45$, and a spin ensemble size of $N=4$ so that $\mathrm{P}_{max}=2$.
A subtle but crucial aspect of cooling the oscillator is to choose the detuning $\epsilon$ precisely so that the conditional displacement operator $D(\theta)$ in Fig. 2 is not perfectly aligned with the position quadrature. Note that when the detuning is zero, $D(\theta)$ is just perfectly aligned with the momentum quadrature. In both cases the oscillator cannot be cooled [58]. This observation is intriguingly related to Gottesman-Kitaev-Preskill (GKP) state encoding by phase estimation, which uses the same control of applying inversion pulses to the probe spin and its measurement. However, the detuning $\epsilon$ must now be carefully chosen for each measurement to ensure that the conditional oscillator displacement is exactly aligned with either the momentum or position quadratures. In addition, we must carefully tune the magnitude of the oscillator displacement. Starting with an oscillator in its zero photon ground state, this process actually increases the number of photons and allows the preparation of GKP states in the oscillator [61, 62].
In the Appendix A, we discuss this proposal in more detail, where it becomes clear that encoding GKP states in a nanomechanical oscillator would be much more challenging. This is because a much stronger spin-mechanical coupling strength is needed, since the coherence time of mechanical vibration is rather limited.
VI Discussion
In this research, we have emphasized the profound influence that the precise control of a single spin can have on the dynamics of ensemble oscillator systems, and demonstrated this effect in the regime of weak coupling. This finding is central to a wide range of quantum technological applications. Specifically, we have introduced a ground-state cooling protocol for a tripartite hybrid quantum system that includes a spin ensemble, an oscillator, and a single probe spin. By exploiting the weak dispersive coupling between the spins and the oscillator, together with the feedback from frequent measurements of the probe spin, our protocol can cool both the spin ensemble and the oscillator to their ground states, either individually or in tandem.
Our numerical simulations over various parameter settings not only confirm the practicality of our cooling protocol, but also provide a basis for its experimental implementation. For example, our analysis shows that a spin ensemble of $N=14$ spins could be cooled with a spin-mechanical coupling strength as low as 1 kHz at a mechanical oscillation frequency of about 1 MHz. Such coupling strengths, though are on the higher side of current state-of-the-art experiments, can become realizable by further reducing dimensions of the mechanical oscillator [63]. For other solid-state defects in silicon carbide, the coupling strengths could reach a few MHz [53], allowing us to attain higher cooling rates.
Another notable limitation of our cooling scheme is its reliance on post-selection, which leads to an exponential decrease in success rate as the number of probe measurements increases. Nevertheless, our protocol offers a distinct advantage in its applicability to cooling large spin ensembles that are otherwise difficult to control and cannot be cooled by conventional optical or energy transfer methods. Furthermore, for spin ensembles that are amenable to such traditional cooling methods, such as NV centers in diamond, our approach adds to the toolbox and serves as a complementary polarization method. For example, optical methods could initially polarize an NV ensemble to a certain degree, after which our protocol could be used to further cool the ensemble toward the ground state, while maintaining a reasonable success rate. This will help us to alleviate the heating problem caused by intensive optical initialization.
Finally, this work also explores the potential of extending the cooling scheme to encode complex quantum states in oscillators, with a particular focus on the GKP states. Despite the important role of these states in quantum error correction [61, 62] and quantum sensing [64], their creation in macroscopic objects, a such as a nano-mechanical oscillator considered in this work, promises to deepen our understanding of the boundary between the classical and quantum worlds [65, 66]. Despite the inherent challenges, in particular the need for much stronger coupling strengths for experimental realization, our results provide insights for future efforts in this direction.
@@ Acknowledgement
The authors would like to thank Ruoming Peng and Jixing Zhang for feedback on the manuscript. This work is supported by German Federal Ministry of Education and Research (BMBF) for the project QECHQS under Grant No. 16KIS1590K.
Appendix A Encoding a qubit into a macroscopic oscillator
In this section, we consider a hybrid system consisting of only the probe spin and the oscillator. We show that we can prepare a macroscopic mechanical oscillator in a complicated quantum state by controlling the probe. The exploration of such quantum superposition states in macroscopic objects has emerged as a significant area of interest, mainly because of its potential to explore the frontier between classical and quantum physics [66, 65].
Our proposal is realized by carefully tuning the detuning ($\epsilon$) and applying a certain number of inversion pulses to the auxiliary spin. The quantum state in which we propose to place a macroscopic mechanical oscillator is the so-called Gottesman-Kitaev-Preskill (GKP) state [61, 67], which are defined as the simultaneous eigenstates of the two stabilizers, i.e.,

### v2

@@ C Tuning dynamics through probe projections
To cool the ensemble-oscillator subsystem, we repeatedly implement the circuit shown in Fig. 2(a). It is important to emphasize that, each time, only a probe measurement result of 0 is accepted, otherwise the whole process is restarted. This process effectively implements the projector $\mathcal{P}_{\scriptscriptstyle N,+}$ given in Eq. (14) many times. The probability that the ensemble is projected into the $\mathrm{I}_{k}$ sector after $M$ repetitions is calculated as
$\mathrm{P}_{m}(M)=\frac{1}{\sum_{m}\mathrm{P}_{m}}\mathrm{Tr}[\mathcal{P}_{\scriptscriptstyle N,+}^{M}\,\rho_{B}(0)\,\mathcal{P}_{\scriptscriptstyle N,+}^{M}],$ (16)
where $m=(2k-N)/2$ is the total magnetization of the ensemble. This dependence on $m$ shows that, with appropriate parameters, it is possible to polarize the ensemble, i.e., to obtain $\mathrm{P}{\scriptscriptstyle N/2}(M)\approx 1$ when $M$ is large enough.
IV System cooling
In this section, we describe how the repeated probe projections can impose a thermal filter on the hybrid system, allowing for controlled cooling of either the oscillator, the ensemble, or both. We focus primarily on the weak coupling regime, where the coupling strength $g$ is significantly smaller than the mechanical oscillation frequency $\omega$, denoted as $g\ll\omega$. To explore this regime, we set the spin coupling strengths $g,~{}g_{0}$ to $10$ kHz, while the mechanical oscillation frequency $\omega$ is chosen to be $1.2$ MHz. Although such a frequency for the mechanical oscillator is experimentally achievable [45, 46], the chosen values for the spin-mechanical couplings are at the upper end of what is typically achievable in practice [43].
These choices represent a compromise between stronger coupling for faster cooling and additional noise due to imperfect pulses on the probe. As we discussed in Eq. (11), at least $\mathcal{O}(\omega/g_{0})$ pulses are required for each probe projection to ensure that the effective spin-oscillator coupling is comparable to the oscillator frequency, which becomes clear in Eq. (9). However, a higher number of pulses would result in a longer total evolution time to implement the cooling protocol, which is undesirable in the presence of decoherence. Throughout this work, we set the number of pulses in each round of probe projection to $n=\omega/g_{0}=120$, as shown in Fig. 1(b).
Figure 3: Pulse Detuning: A key parameter in the cooling scheme. (a) This panel shows the intricate dependence of achieving complete spin polarization on pulse detuning ($\epsilon$). The probability that all ensemble spins are aligned upward after $M=100$ probe projections, denoted by $\mathrm{P}_{\scriptscriptstyle m=N/2}(M)$ and derived in Eq. (16), is plotted on the vertical axis. This probability is affected by both the number of spins in the ensemble and the size of $\epsilon$, with $\epsilon=2$ kHz already appearing too large for effective polarization of the ensemble. (b) If we choose an even larger detuning ($\epsilon=2.5$ kHz) and set the number of spins to $N=4$, we observe a rapid decrease of the thermal occupancy of the oscillator to zero, indicating that the oscillator is cooling. However, the spin polarization remains almost unchanged and reaches a steady state of 0.0625 after about 40 probe projections. Note that we have considered an ensemble of $N=4$ spins, so the y-axis represents the relative value of $\mathrm{P}_{z}(M)/2$. Correspondingly, the y-axis also represents the relative value of the thermal occupancy of the oscillator with respect to its initial value, which is set to about 45. Further simulation details are summarized in Appendix A.
A Pulse detuning as a control parameter
The probability $\mathrm{P}_{m}(M)$ in Eq. (16) depends non-trivially on the filter function in Eq. (13) that is imposed by pulses on the probe. As discussed in our previous work [43], varying the value of $\epsilon$ can lead to cooling, heating, or squeezing of the oscillator. In particular, Ref. [43] has found that a ratio of $\epsilon/\omega\approx 5\times 10^{-3}$ leads to rapid cooling of the oscillator, which corresponds to $\epsilon\approx 6$kHz for the oscillator frequency of $1.2$ MHz.
However, to achieve complete polarization of the ensemble, we must avoid premature cooling of the oscillator before the ensemble is polarized. This is due to the significant effect of the thermal occupancy of the oscillator on the indirect interaction strength between the probe and the ensemble, as elaborated in Eq. (1). In particular, at zero thermal occupancy, the oscillator-mediated effective interaction between the probe and the ensemble spins goes correspondingly to zero.
First, we run simulations with the pulse detuning $\epsilon$ close to zero to explore the near-resonance regime. By adjusting its magnitude from 0 to 2 kHz and varying the number of spins in the ensemble, our simulations reveal a nuanced relationship between the probability of achieving complete ensemble polarization and the pulse detuning $\epsilon$, as shown in Fig. 3(a). To cool the spin ensemble, the choice of $\epsilon$ should satisfy the condition that the probe-oscillator coupling is enhanced without prematurely cooling the oscillator. Consequently, satisfying this condition requires fine-tuning of all parameters of the Hamiltonian. This makes it difficult to determine the exact dependence of the optimal pulse detuning $\epsilon$ on the number of ensemble spins $N$ and the number of probe projections $M$, as shown in Fig. 3(a). For example, it can be seen that within an ensemble of constant size, varying the detuning can lead to very different results when implementing the proposed cooling protocol. Another important observation from Fig. 3(a) suggests that a detuning of $\epsilon=2$ kHz may be large enough to facilitate rapid oscillator cooling and thereby prevent ensemble polarization.
To investigate this further, we introduce an even larger detuning of $\epsilon=2.5$ kHz for an ensemble of $N=4$ spins. The simulation results shown in Fig. 3(b) indicate that at this substantial pulse detuning, the oscillator quickly reaches its ground state after about 80 probe projections. This premature cooling of the oscillator reduces the interaction between the probe and the ensemble, so that the ensemble remains largely unpolarized. The ensemble polarization fluctuates only slightly around the final saturation value for the first few tens of probe projections, as shown in the inset of Fig. 3(b).
Furthermore, as an example, we search for the optimal pulse detuning for cooling a larger ensemble of $N=14$ spins. An exhaustive parameter search leads us to set the detuning to $\epsilon=1.16$ kHz. This setting shows that the sector with all spins pointing up (i.e. $m=7$) is predominantly left after $M=100$ probe projections. The probability of full polarization is as high as 95.82%, as shown in Fig. 2(b).
The above results underscore the critical role of $\epsilon$ as a controlling parameter in the cooling protocol. In addition, a subtle and interesting aspect of cooling the oscillator is to choose the detuning $\epsilon$ precisely so that the conditional oscillator displacement operator is not perfectly aligned with the position or momentum quadrature. Otherwise, the oscillator will be squeezed instead of cooled [43]. Note that, in the resonant case that $\epsilon=0$, the displacement operator of the oscillator is along the position quadrature.
This observation is interestingly related to the encoding of Gottesman-Kitaev-Preskill (GKP) states, which have applications in quantum sensing and quantum error correction [47, 48, 49]. Its encoding can be realized by the same control of the probe spin as our cooling protocol. Starting with an oscillator in its zero photon ground state, repeated probe projections can actually increase the number of photons and leave the oscillator in GKP states [47, 48]. The generation of these states in macroscopic objects could enrich our understanding of the quantum-classical interface, notwithstanding their important roles in quantum error correction [47, 48] and quantum sensing [50]. More details of GKP encoding can be found in Appendix C.
@@ A Pulse detuning as a control parameter
Figure 4: Cooling of both the oscillator and the ensemble. This figure shows the polarization $\mathrm{P}_{z}(M)$ and the thermal occupancy $n(\omega)$ for both the spin ensemble and the oscillator, illustrating how the cooling/polarization proceeds with the number of probe projections $M$. The effective cooling of these quantum systems occurs in different parameter regimes. Initially, the spin ensemble is cooled by adjusting the pulse intervals to near-resonance conditions, specifically $\tau_{1}$ such that $\epsilon_{1}=1.1$ kHz. In this case, the thermal occupancy of the oscillator remains largely unaffected. After polarizing the spin ensemble, we change the pulse interval to $\tau_{2}$ with a pulse detuning of $\epsilon_{2}=5$ kHz. This more significant detuning leads to a rapid cooling of the oscillator to its ground state. Note that the number of ensemble spins is $N=4$ so that the y-axis represents the relative value of spin polarization $\mathrm{P}_{z}(M)/2$. The oscillator thermal occupancy is also plotted as its relative to the initial occupancy $n_{0}\approx 45$. Further simulation details are summarized in Appendix A.
B Simultaneous cooling of ensemble and oscillator
In this section, we discuss the cooling of both the spin ensemble and the oscillator simultaneously. The key to achieving this goal is to avoid premature cooling of the oscillator, which should instead occur after the ensemble is fully polarized. This requires tuning the pulse detuning $\epsilon$ separately for the cooling phases of the oscillator and the ensemble. Fig. 4 illustrates a pulse sequence that effectively polarizes an ensemble of $N=4$ spins while cooling an oscillator from an initial thermal occupancy of $n_{0}\approx 45$.
During this simulation, we find that the optimal detuning that achieves complete polarization of the ensemble is identified as $\epsilon_{1}=1.1$kHz, while the oscillator occupancy is not much reduced. In this case, performing 250 probe projections can achieve complete polarization of the ensemble. Then, to cool the oscillator, we adjust the detuning to a higher value, $\epsilon_{2}=5$kHz, and execute an additional series of $50$ probe projections. This transition in the magnitude of the pulse detuning results in a rapid cooling of the oscillator, as shown in the right part of Fig. 4.
It is important to note that increasing $\epsilon_{2}$ requires a higher number of inversion pulses on the probe to achieve a comparable cooling effect, as can be seen from Eq. (13). However, implementing a higher number of inversion pulses presents experimental challenges, primarily due to the potential for pulse imperfections. This consideration emphasizes the need for precise control and optimization in the experimental setup.
V Discussion
In this study, we have developed a novel ground-state cooling protocol for a tripartite hybrid quantum system consisting of a spin ensemble, an oscillator, and a probe spin. Central to this protocol is the exploitation of the feedback provided by frequent measurements of the prob spin. Our analysis has shown that even with weak dispersive coupling between the spins and the oscillator, one can effectively cool both the ensemble and the oscillator to their ground states, either separately or simultaneously. We would also like to emphasize that, except for the probe, there is no need for quantum coherence in other parts of the hybrid system, marking a clear distinction from resonant transfer schemes. The polarization considered here can also be viewed as a remote sensing scheme when considering a macroscopic oscillator, where the boundaries of the sensing volume can be extended beyond the limits set by phase accumulation schemes [25].
Our numerical simulations, performed over a range of parameters, underscore the feasibility of the cooling protocol, and lay the groundwork for its experimental realization. For example, our results indicate that a spin ensemble consisting of $N=14$ spins could be cooled with a spin-mechanical coupling strength of only a few kHz at a mechanical oscillation frequency close to 1 MHz. Although such coupling strengths are within the experimental reach, the cooling efficiency can be further improved by miniaturization of the mechanical oscillator [51]. In the context of solid-state defects in silicon carbide, coupling strengths can reach a few MHz [37], promising even more efficient cooling rates. In addition, the use of magnons instead of NMOs could significantly enhance the spin-oscillator coupling, facilitating the experimental adoption of our protocol.
A notable challenge of our cooling approach is its dependence on post-selection, which results in a decreasing success rate with increasing number of probe projections. Nevertheless, our protocol offers a significant advantage for cooling large spin ensembles that are traditionally difficult to handle and cannot be cooled by standard optical or energy transfer techniques. For spin ensembles that are amenable to conventional cooling methods, such as NV centers in diamond, our protocol adds to the cooling toolbox.
This research highlights the significant impact that precise control of a single spin can have on the dynamics of a hybrid system, enabling energy transfer-like effects beyond the inherent capabilities of weak dispersive coupling. In addition to remote sensing, these results also provide insight into the preparation of complex quantum states such as the GKP states in a macroscopic object. Furthermore, the ability to interact/communicate with a few target spins within a large network may also be useful in quantum network protocols where a similar result is desired.
@@ Acknowledgement
The authors would like to thank Jixing Zhang, Xin Zhang, and Ruoming Peng for their valuable feedback on the manuscript. This work was supported by the German Federal Ministry of Education and Research (BMBF) through the project QECHQS with Grant No. 16KIS1590K, and through the Future Cluster QSens. Additional support was provided by the Baden-Württemberg Stiftung through the project SPOC with Grant agreement No. QT-6. The Deutsche Forschungsgemeinschaft (DFG) has funded this work under GRK 2642, and under Project No. 384846402 as part of FOR 2724.
Appendix A Simulation details
The coupling parameters are also $g_{0}=g=10$ kHz, with an oscillator frequency of $\omega=1.2$ MHz. Accordingly, we set the number of inversion pulses for each round of probe projection to $n=120$, see the probe control sequence in Fig. 11(b). This is to ensure that the effective probe-oscillator coupling is comparable to the oscillator frequency, as discussed in Eq. (11).
In this work, we perform the simulation by calculating the dynamics directly on the initial Hamiltonian given in Eq. (1) without any approximation. In addition, for simplicity, we approximate the oscillator mode as having only 100 levels, which is expected to be sufficient since we assume an initial oscillator thermal occupation of $n_{0}\approx 45$ in the simulations.
Appendix B Effect of decoherence
Our numerical simulations have not considered the direct effect of spin dephasing noise. For the hybrid system considered in this work, such noise is induced by the spin bath surrounding the probe and the ensemble, and is expected to be the dominant noise source for the hybrid system considered [52]. In each measurement, both the ensemble spins and the spin bath surrounding them collapse partially and simultaneously towards the steady states. This collapse is sensitive only to the probe population in the measurement basis [53].
Similarly, in the case of the oscillator, its dissipation ($T_{1}$) also has a negligible effect. As shown in Fig. 4, the cooling of the spin ensemble requires more measurements and time compared to the cooling of the oscillator. Since the (initial) state of the oscillator is unaffected during the cooling of the ensemble, the effect of the oscillator’s $T_{1}$ can be ignored. Counterintuitively, in some cases rapid decoherence of the intra-bath correlations can actually increase the cooling rate rather than decrease it [54]. See further discussion of the effect of oscillator $T_{1}$ decoherence (relaxation) on its cooling in our previous work [43].
Appendix C Encoding a qubit into a macroscopic oscillator
In this section, we consider a hybrid system consisting only of the probe spin and the oscillator. We will show that we can prepare a macroscopic mechanical oscillator into a complicated quantum state using the same probe control sequence as our cooling protocol in the main text. The exploration of such quantum superposition states in macroscopic objects has emerged as a significant area of interest, mainly because of its potential to explore the frontier between classical and quantum physics [55, 56].
The quantum state in which we propose to prepare into a macroscopic mechanical oscillator is the so-called Gottesman-Kitaev-Preskill (GKP) state [47, 57], which is defined as the simultaneous eigenstates of the two stabilizers, i.e.,